# Rejuvenating Cross-Entropy Loss in Knowledge Distillation for Recommender Systems

### v1

@@ Abstract
This paper analyzes Cross-Entropy (CE) loss in knowledge distillation (KD) for recommender systems. KD for recommender systems targets at distilling rankings, especially among items most likely to be preferred, and can only be computed on a small subset of items. Considering these features, we reveal the connection between CE loss and NDCG in the field of KD. We prove that when performing KD on an item subset, minimizing CE loss maximizes the lower bound of NDCG, only if an assumption of closure is satisfied. It requires that the item subset consists of the student’s top items. However, this contradicts our goal of distilling rankings of the teacher’s top items. We empirically demonstrate the vast gap between these two kinds of top items. To bridge the gap between our goal and theoretical support, we propose Rejuvenated Cross-Entropy for Knowledge Distillation (RCE-KD). It splits the top items given by the teacher into two subsets based on whether or not it is ranked highly by the student. For the subset that defies the condition, a sampling strategy is devised to use teacher-student collaboration to approximate our assumption of closure. We also combine the losses on the two subsets adaptively. Extensive experiments demonstrate the effectiveness of our method. Our code is available at https://anonymous.4open.science/r/RCE-KD.
@@ 1 Introduction
Recently, with the scaling law in recommender systems Zhai et al. (2024) being gradually discovered, many researchers have proposed extremely large models Ohsaka & Togashi (2023); Zhai et al. (2024) to pursue better recommendation accuracy. However, the increase in model size inevitably incurs high storage costs and inference latency, causing higher maintenance costs and lower user satisfaction.
To improve the inference efficiency and decrease the storage cost of recommendation models without sacrificing their recommendation accuracy, knowledge distillation (KD) for recommender systems Kang et al. (2020); Sun et al. (2024) has attracted attention. KD Hinton et al. (2015) is an approach for model compression. It aims to transfer knowledge from a pre-trained large teacher to a small student. Once training is complete, only the small student is used for inference. Among existing works on KD, response-based KD Hinton et al. (2015) encourages students to mimic the teacher’s predictions and has gained extreme attention due to its excellent performance. As a popular loss for response-based KD methods, Cross-Entropy (CE) loss is very important. Most response-based KD methods Huang et al. (2022); Cui et al. (2023) in Computer Vision (CV) and Natural Language Processing (NLP) are based on CE loss. However, little work has been done to use or analyze CE loss in KD for recommender systems. Note that KD for recommender systems has two unique features: 1) It focuses more on rankings than specific scores, especially among the teacher’s top items Kang et al. (2020). 2) KD can only be conducted on a small subset of items since the quantity of all the items is very large. These features make the compatibility of CE loss in KD for recommender systems questionable. To obtain an initial insight into the performance of CE loss, we present the results of vanilla CE loss and several response-based KD methods in Figure 1. To cover as many types of loss functions as possible, we consider the point-wise loss (i.e., CD Lee et al. (2019)), pair-wise loss (i.e., UnKD Chen et al. (2023)), and RRD-based losses Kang et al. (2020) (a list-wise loss, i.e., RRD Kang et al. (2020) and HetComp Kang et al. (2023)). In vanilla CE loss, we compute CE loss using the teacher’s top items. We find that vanilla CE loss is often inferior to all baselines. This result contrasts with the extensive use of CE loss for KD in other fields.
@@
However, full-item KD is not practical due to the extremely large number of items. In real-world scenarios, CE loss could only be computed on a subset of items (i.e., the partial-item KD), such as the teacher’s top items in vanilla CE. For this case, we define partial NDCG, which only considers rankings within a subset of items. Then, we prove that CE loss bounds partial NDCG. However, it holds only if the item subset satisfies our assumption of closure (Assumption 4.3). It requires that all items that the student ranks higher than any item in the subset are also in the subset. This assumption emphasizes the effect of the student’s top items. Recall that our goal is to distill rankings among the teacher’s top items Reddi et al. (2021). Unfortunately, we observe that the top items given by the teacher are usually ranked low by the student. This makes it difficult for the teacher’s top items to satisfy our assumption of closure. Thus, vanilla CE cannot bound partial NDCG and performs poorly.
@@ 2.1 Knowledge Distillation for Recommender Systems
Response-based methods focus on teachers’ predictions. CD Lee et al. (2019) samples unobserved items from a distribution associated with their rankings predicted by students, and distills with a point-wise loss. RankDistil Reddi et al. (2021) enables students to mimic teachers by sampling high-ranking items predicted by teachers and calculating multiple forms of loss functions on them. RRD Kang et al. (2020) adopts a list-wise loss to maximize the likelihood of the teacher’s recommendation list. Note that RRD could be regarded as the extension of ListMLE loss Xia et al. (2008) to the top-K setting Xia et al. (2009). Based on RRD, DCD Lee & Kim (2021) uses the discrepancy between the teacher and student model predictions to decide which knowledge to distill. HetComp Kang et al. (2023) transfers the ensemble knowledge of heterogeneous teachers by constructing easy-to-hard knowledge sequences from the teachers’ trajectories.
Feature-based methods focus on the intermediate representations of the teacher. FreqD Zhu & Zhang (2024) defines knowledge as different frequency components of the features and proposes emphasizing important knowledge by graph filtering. PCKD Zhu & Zhang (2025) observes that projectors in feature-based KD interrupt user preference contained in the features and designs two regularization terms to restrict the projectors.
Relation-based methods focus on the relationships between different items. HTD Kang et al. (2021) distills the sample relation hierarchically to alleviate the capacity gap between the student and teacher.
@@ 2.2 Connection between CE Loss and NDCG
Recently, many studies Cao et al. (2007); Ravikumar et al. (2011); Bruch et al. (2019); Wu et al. (2024); Yang et al. (2024); Xu et al. (2024) on learning-to-rank (LTR) have focused on the impact of different surrogate loss functions on NDCG. Among them, CE loss is of particular interest due to its wide range of applications. As a pioneer, ListNet Cao et al. (2007) introduces CE loss into LTR by defining the top-one probability. Then, Bruch et al. (2019) for the first time proves that CE loss is a bound on NDCG when considering binary ground-truth labels. Subsequently, work has been done to improve CE loss based on this conclusion. For example, PSL Yang et al. (2024) changes the surrogate activations, and SCE Xu et al. (2024) increases the weight of negative samples in CE loss to achieve a tighter bound of NDCG. Another work relevant to us is Wu et al. (2024). It reveals the pros and cons of sampled CE loss for item recommendation and also relates it to NDCG. However, these methods mentioned above hardly address the case of non-binary ground-truth labels. Moreover, they either do not focus on the scenarios that need item sampling or simply use uniform sampling without making any assumptions about the items being sampled. This makes them entirely inapplicable for KD, where we take the teacher’s predictions as labels and emphasize the top-ranked items.
@@ 3.2 Cross-Entropy Loss for Knowledge Distillation
where $r_{ui}^{T}$ and $r_{ui}^{S}$ denote the scores predicted by the teacher and the student, respectively. $\sigma(r_{ui}^{T},\mathcal{J}^{u})=\exp(r_{ui}^{T})/\sum_{j\in\mathcal{J}^{u}}\exp(r_{uj}^{T})$ denotes the softmax over item set $\mathcal{J}^{u}$. Similarly $\sigma(r_{ui}^{S},\mathcal{J}^{u})=\exp(r_{ui}^{S})/\sum_{j\in\mathcal{J}^{u}}\exp(r_{uj}^{S})$. Note that for each user $u$, we only have access to a sampled item subset since it is computationally intractable over the entire item set $\mathcal{I}$ Sun et al. (2024).
@@ 4.3 Challenge in Partial-Item KD
According to our analysis, $\mathcal{J}^{u}$ must satisfy Assumption 4.3 for the connection between CE loss and partial NDCG to hold. However, it is difficult to satisfy this assumption if we do not explicitly consider the student’s top items. Specifically, in Figure 2, we report the relationship between the student’s and the teacher’s rankings at the beginning and end of the training. The student is trained with vanilla CE loss, which is computed using the teacher’s top items. The dataset is CiteULike. Detailed analysis and results on all datasets are given in Appendix A. From the results, we find that:
@@ 5.1 Overview of RCE-KD
To unleash the potential of CE loss of distilling rankings among the teacher’s top items, we propose RCE-KD, a novel approach involving both the teacher’s and the student’s top items in KD. The key is to split the teacher’s top items into two subsets based on whether or not an item is ranked high by the student. Then, we try to make both item subset satisfy Assumption 4.3 exactly or approximately.
Let $\mathcal{Q}^{T}_{u}\triangleq\mathop{\text{arg top$K$}}(\bm{r}_{u}^{T})$ and $\mathcal{Q}^{S}_{u}\triangleq\mathop{\text{arg top$K$}}(\bm{r}_{u}^{S})$ denote the sets of top-$K$ items predicted by the teacher and the student, respectively. We aim to transfer the teacher’s ranking ability over $\mathcal{Q}_{u}^{T}$ Kang et al. (2020); Lee et al. (2019); Tang & Wang (2018). In RCE-KD, we propose to separate $\mathcal{Q}_{u}^{T}$ into two subsets. The first subset is the interaction between $\mathcal{Q}_{u}^{T}$ and $\mathcal{Q}_{u}^{S}$. The second subset contains the remaining items in $\mathcal{Q}_{u}^{T}$. Formally, $(\mathcal{Q}_{u}^{T})_{1}\triangleq\mathcal{Q}_{u}^{T}\cap\mathcal{Q}_{u}^{S}\quad\text{and}\quad(\mathcal{Q}_{u}^{T})_{2}\triangleq\mathcal{Q}_{u}^{T}\backslash(\mathcal{Q}_{u}^{T})_{1}.$
@@ 5.4 Adaptive Loss Fusion
Note that $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ play different roles. $\mathcal{L}_{1}$ focuses on $(\mathcal{Q}_{u}^{T})_{1}$, consisting of top items considered by both the student and teacher. The goal on these items is to distill their fine-grained rankings, which is done by exactly maximize the partial NDCG. On the contrary, $\mathcal{L}_{2}$ focuses on items not well-mastered by the student. The goal for these items is to improve their rankings, which is done by making the student imitate the teacher’s ranking on $(\mathcal{Q}_{u}^{T})_{2}$ and randomly sampled items.
@@ 6 Experiments
Section 6.1 first introduces the experimental settings. The implementation details are shown in Appendix C.1. Then, the overall performance comparison is shown in Section 6.2. Consequently, we investigate the training efficiency of all compared KD methods in Section 6.3. The ablation study is conducted in Section 6.4. To verify our sampling strategy’s efficiency for approximating Assumption 4.3, we conduct experiments in Appendix C.2. We present hyperparameter analysis in Appendix C.4. Finally, in Appendix C.5, we demonstrate the effectiveness of applying our method to sequential recommendation to showcase its generalization capability for recommendation tasks.
@@ 6.1 Experimental Settings
Datasets. We conduct experiments on three public datasets, including CiteULike Wang et al. (2013); Kang et al. (2022; 2021), Gowalla Cho et al. (2011); Tang & Wang (2018); Lee et al. (2019), and Yelp2018 Lee et al. (2019); Kweon et al. (2021). Detailed statistics and methods of constructing training and test sets are given in Appendix C.1.
@@
Baselines. We compare our method with five response-based KD methods: CD Lee et al. (2019), RRD Kang et al. (2020), DCD Lee & Kim (2021), and HetComp Kang et al. (2023). The introduction of these methods is in Appendix C.1.
Backbones. We refer to previous works Chen et al. (2023); Kang et al. (2020; 2021), and use MF Rendle et al. (2012) and LightGCN He et al. (2020). We also add HSTU Zhai et al. (2024) as a new backbone, which is a popular generative recommendation model.
@@ 6.3 Training Efficiency
In this section, we report the training efficiency of our method and comparison methods. All results are obtained by testing with PyTorch on a GeForce RTX 3090 GPU.
@@ 6.4 Ablation Study
RCE-KD consists of three key components: 1) It divides the teacher’s top items into two subsets; 2) It computes CE loss on items selected from both the teacher’s and the student’s top items; 3) An adaptive mechanism is proposed to combine the losses on these subsets. To validate the effectiveness of these key components, we design four variants: 1) RCE-KD w/o sep does not compute losses for $(\mathcal{Q}_{u}^{T})_{1}$ and $(\mathcal{Q}_{u}^{T})_{1}$, separately. It only computes CE loss on $\mathcal{A}^{u}\cup\mathcal{Q}_{u}^{S}$; 2) RCE-KD w/o S only aligns the predictions on the teacher’s predicted top items, i.e., $\mathcal{Q}_{u}^{T}$; 3) Similarly, RCE-KD w/o T only aligns the predictions on the student’s predicted top items, i.e., $\mathcal{Q}_{u}^{S}$; 4) RCE-KD w/ const replaces the adaptive weight derived from Eq.(10) with a constant hyperparameter $\gamma$.
Figure 3 shows the results of these four variants on Gowalla and Yelp, and three Teacher/Student settings. The results of the remaining settings are provided in Appendix C.3. We find that all variants are inferior to the original RCE-KD, which demonstrates the effectiveness of all key components. Moreover, RCE-KD w/o S usually performs worse than RCE-KD w/o T. We believe that the reason is that the top items given by the student can exactly satisfy Assumption 4.3, while the top items given by the teacher do not. The superiority of RCE-KD w/ const over RCE-KD w/o T demonstrates the necessity of involving top items from both the student and the teacher. Finally, the superiority of RCE-KD over RCE-KD w/ const and RCE-KD w/o sep validates the effectiveness of our adaptive weighting scheme and the necessity of splitting out the two subsets and treating them separately.

### v2

# arXiv:2509.20989v2[cs.IR] 01 Mar 2026

@@ Abstract
This paper analyzes Cross-Entropy (CE) loss in knowledge distillation (KD) for recommender systems. KD for recommender systems targets at distilling rankings, especially among items most likely to be preferred, and can only be computed on a small subset of items. Considering these features, we reveal the connection between CE loss and NDCG in the field of KD. We prove that when performing KD on an item subset, minimizing CE loss maximizes the lower bound of NDCG, only if an assumption of closure is satisfied. It requires that the item subset consists of the student’s top items. However, this contradicts our goal of distilling rankings of the teacher’s top items. We empirically demonstrate the vast gap between these two kinds of top items. To bridge the gap between our goal and theoretical support, we propose Rejuvenated Cross-Entropy for Knowledge Distillation (RCE-KD). It splits the top items given by the teacher into two subsets based on whether they are highly ranked by the student. For the subset that defies the condition, a sampling strategy is devised to use teacher-student collaboration to approximate our assumption of closure. We also combine the losses on the two subsets adaptively. Extensive experiments demonstrate the effectiveness of our method. Our code is available at https://github.com/BDML-lab/RCE-KD.
@@ 1 Introduction
Recently, as the scaling law in recommender systems (Zhai et al., 2024) has been increasingly recognized, many researchers have proposed extremely large models (Ohsaka and Togashi, 2023; Zhai et al., 2024) to achieve higher recommendation accuracy. However, the increase in model size inevitably incurs high storage costs and inference latency, causing higher maintenance costs and lower user satisfaction.
To improve the inference efficiency and decrease the storage cost of recommendation models without sacrificing their recommendation accuracy, knowledge distillation (KD) for recommender systems (Kang et al., 2020; Sun et al., 2024) has attracted attention. KD (Hinton et al., 2015) is an approach for model compression. It aims to transfer knowledge from a pre-trained large teacher to a small student. Once training is complete, only the small student is used for inference. Among existing works on KD, response-based KD (Hinton et al., 2015) encourages students to mimic the teacher’s predictions and has gained extreme attention due to its excellent performance. As a popular loss for response-based KD methods, Cross-Entropy (CE) loss is very important. Most response-based KD methods (Huang et al., 2022; Cui et al., 2023) in Computer Vision (CV) and Natural Language Processing (NLP) are based on CE loss. However, little work has been done to use or analyze CE loss in KD for recommender systems. Note that KD for recommender systems has two unique features: 1) It focuses more on rankings than specific scores, especially among the teacher’s top items (Kang et al., 2020). 2) KD can only be conducted on a small subset of items since the quantity of all the items is very large. These features make the compatibility of CE loss in KD for recommender systems questionable. To obtain an initial insight into the performance of CE loss, we present the results of vanilla CE loss and several response-based KD methods in Figure 1. To cover as many types of loss functions as possible, we consider the point-wise loss (i.e., CD (Lee et al., 2019)), pair-wise loss (i.e., UnKD (Chen et al., 2023)), and RRD-based losses (Kang et al., 2020) (a list-wise loss, i.e., RRD (Kang et al., 2020), HetComp (Kang et al., 2023)), and our method RCE-KD. In vanilla CE loss, we compute CE loss using the teacher’s top items. We find that vanilla CE loss is often inferior to all baselines. This result contrasts with the extensive use of CE loss for KD in other fields.
@@
However, full-item KD is not practical due to the extremely large number of items. In real-world scenarios, CE loss could only be computed on a subset of items (i.e., the partial-item KD), such as the teacher’s top items in vanilla CE. For this case, we define partial NDCG, which only considers rankings within a subset of items. Then, we prove that CE loss bounds partial NDCG. However, it holds only if the item subset satisfies our assumption of closure (Assumption 4.3). It requires that all items that the student ranks higher than any item in the subset are also in the subset. This assumption emphasizes the effect of the student’s top items. Recall that our goal is to distill rankings among the teacher’s top items (Reddi et al., 2021). Unfortunately, we observe that the top items given by the teacher are usually ranked low by the student. This makes it difficult for the teacher’s top items to satisfy our assumption of closure. Thus, vanilla CE cannot bound partial NDCG and performs poorly.
@@ 2.1 Knowledge Distillation for Recommender Systems
Response-based methods focus on teachers’ predictions. CD (Lee et al., 2019) samples unobserved items from a distribution associated with their rankings predicted by students, and distills with a point-wise loss. RankDistil (Reddi et al., 2021) enables students to mimic teachers by sampling high-ranking items predicted by teachers and calculating multiple forms of loss functions on them. RRD (Kang et al., 2020) adopts a list-wise loss to maximize the likelihood of the teacher’s recommendation list. Note that RRD could be regarded as the extension of ListMLE loss (Xia et al., 2008) to the top-K setting (Xia et al., 2009). Based on RRD, DCD (Lee and Kim, 2021) uses the discrepancy between the teacher and student model predictions to decide which knowledge to distill. HetComp (Kang et al., 2023) transfers the ensemble knowledge of heterogeneous teachers by constructing easy-to-hard knowledge sequences from the teachers’ trajectories.
Feature-based methods focus on the intermediate representations of the teacher. FreqD (Zhu and Zhang, 2024) defines knowledge as different frequency components of the features and proposes emphasizing important knowledge by graph filtering. PCKD (Zhu and Zhang, 2025) observes that projectors in feature-based KD interrupt user preference contained in the features and designs two regularization terms to restrict the projectors.
Relation-based methods focus on the relationships between different items. HTD (Kang et al., 2021) distills the sample relation hierarchically to alleviate the capacity gap between the student and teacher.
@@ 2.2 Connection between CE Loss and NDCG
Recently, many studies (Cao et al., 2007; Ravikumar et al., 2011; Bruch et al., 2019; Wu et al., 2024; Yang et al., 2024; Xu et al., 2024b) on learning-to-rank (LTR) have focused on the impact of different surrogate loss functions on NDCG. Among them, CE loss is of particular interest due to its wide range of applications. As a pioneer, ListNet (Cao et al., 2007) introduces CE loss into LTR by defining the top-one probability. Then, (Bruch et al., 2019) for the first time proves that CE loss is a bound on NDCG when considering binary ground-truth labels. Subsequently, work has been done to improve CE loss based on this conclusion. For example, PSL (Yang et al., 2024) changes the surrogate activations, and SCE (Xu et al., 2024b) increases the weight of negative samples in CE loss to achieve a tighter bound of NDCG. Another work relevant to us is (Wu et al., 2024). It reveals the pros and cons of sampled CE loss for item recommendation and also relates it to NDCG. However, these methods mentioned above hardly address the case of non-binary ground-truth labels. Moreover, they either do not focus on the scenarios that need item sampling or simply use uniform sampling without making any assumptions about the items being sampled. This makes them entirely inapplicable for KD, where we take the teacher’s predictions as labels and emphasize the top-ranked items.
@@ 3.2 Cross-Entropy Loss for Knowledge Distillation
where $r_{ui}^{T}$ and $r_{ui}^{S}$ denote the scores predicted by the teacher and the student, respectively. $\sigma(r_{ui}^{T},\mathcal{J}^{u})=\exp(r_{ui}^{T})/\sum_{j\in\mathcal{J}^{u}}\exp(r_{uj}^{T})$ denotes the softmax over item set $\mathcal{J}^{u}$. Similarly $\sigma(r_{ui}^{S},\mathcal{J}^{u})=\exp(r_{ui}^{S})/\sum_{j\in\mathcal{J}^{u}}\exp(r_{uj}^{S})$. Note that for each user $u$, we only have access to a sampled item subset since it is computationally intractable over the entire item set $\mathcal{I}$ (Sun et al., 2024).
@@ 4.3 Challenge in Partial-Item KD
According to our analysis, $\mathcal{J}^{u}$ must satisfy Assumption 4.3 for the connection between CE loss and partial NDCG to hold. However, it is difficult to satisfy this assumption if we do not explicitly consider the student’s top items. Specifically, in Figure 2, we report the relationship between the student’s and the teacher’s rankings at the beginning and end of the training. The student is trained with vanilla CE loss, which is computed using the teacher’s top items. The dataset is CiteULike. Detailed analysis and results on all datasets are given in Appendix A.1. From the results, we find that:
@@ 5.1 Overview of RCE-KD
To unleash the potential of CE loss of distilling rankings among the teacher’s top items, we propose RCE-KD, a novel approach involving both the teacher’s and the student’s top items in KD. The key is to split the teacher’s top items into two subsets based on whether or not an item is ranked high by the student. Then, we try to make both item subsets satisfy Assumption 4.3 exactly or approximately.
Let $\mathcal{Q}^{T}_{u}\triangleq\mathop{\text{arg top$K$}}(\bm{r}_{u}^{T})$ and $\mathcal{Q}^{S}_{u}\triangleq\mathop{\text{arg top$K$}}(\bm{r}_{u}^{S})$ denote the sets of top-$K$ items predicted by the teacher and the student, respectively. We aim to transfer the teacher’s ranking ability over $\mathcal{Q}_{u}^{T}$ (Kang et al., 2020; Lee et al., 2019; Tang and Wang, 2018). In RCE-KD, we propose to separate $\mathcal{Q}_{u}^{T}$ into two subsets. The first subset is the intersection between $\mathcal{Q}_{u}^{T}$ and $\mathcal{Q}_{u}^{S}$. The second subset contains the remaining items in $\mathcal{Q}_{u}^{T}$. Formally, $(\mathcal{Q}_{u}^{T})_{1}\triangleq\mathcal{Q}_{u}^{T}\cap\mathcal{Q}_{u}^{S}\quad\text{and}\quad(\mathcal{Q}_{u}^{T})_{2}\triangleq\mathcal{Q}_{u}^{T}\backslash(\mathcal{Q}_{u}^{T})_{1}.$
@@ 5.4 Adaptive Loss Fusion
Note that $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ play different roles. $\mathcal{L}_{1}$ focuses on $(\mathcal{Q}_{u}^{T})_{1}$, consisting of top items considered by both the student and teacher. The goal of these items is to distill their fine-grained rankings, which is done by exactly maximize the partial NDCG. On the contrary, $\mathcal{L}_{2}$ focuses on items not well-mastered by the student. The goal for these items is to improve their rankings, which is done by making the student imitate the teacher’s ranking on $(\mathcal{Q}_{u}^{T})_{2}$ and randomly sampled items.
@@
In Appendix C.5, we provide extensive, multi-level experiments that thoroughly validate the superiority of our adaptive $\gamma$-scheduling strategy.
@@ 6 Experiments
Section 6.1 first introduces the experimental settings. The implementation details are shown in Appendix C.1. Then, the overall performance comparison is shown in Section 6.2. Consequently, we investigate the training efficiency of all compared KD methods in Section 6.3. The ablation study is conducted in Section 6.4. To verify our sampling strategy’s efficiency for approximating Assumption 4.3, we conduct experiments in Appendix C.2. We present hyperparameter analysis in Appendix C.4. Then, we provide multi-level experiments in Appendix C.5 to validate the superiority of our adaptive $\gamma$-scheduling strategy. Moreover, in Appendix D.1 and Appendix D.2, we demonstrate the effectiveness of applying our method to sequential recommendation and multi-modal recommendation, respectively, to showcase its generalization capability for recommendation tasks. Finally, we also visualized the evolution of NDCG during training in Appendix A.2 to demonstrate that RCE-KD successfully bounds NDCG as theoretically expected.
@@ 6.1 Experimental Settings
Datasets. We conduct experiments on three public datasets, including CiteULike (Wang et al., 2013; Kang et al., 2022; 2021), Gowalla (Cho et al., 2011; Tang and Wang, 2018; Lee et al., 2019), and Yelp2018 (Lee et al., 2019; Kweon et al., 2021). Detailed statistics and methods of constructing training and test sets are given in Appendix C.1.
@@
Baselines. We compare our method with five response-based KD methods: CD (Lee et al., 2019), RRD (Kang et al., 2020), DCD (Lee and Kim, 2021), HetComp (Kang et al., 2023), and TARec (Zhuang et al., 2025). The introduction of these methods is in Appendix C.1.
Backbones. We refer to previous works (Chen et al., 2023; Kang et al., 2020; 2021), and use MF (Rendle et al., 2012) and LightGCN (He et al., 2020). We also add HSTU (Zhai et al., 2024) as a new backbone, which is a popular generative recommendation model.
@@ 6.3 Training Efficiency
In this section, we report the training efficiency of our method and comparison methods. Since TARec involves a two-stage training process, we did not include it in the comparison. All results are obtained by testing with PyTorch on a GeForce RTX 3090 GPU.
@@ 6.4 Ablation Study
RCE-KD consists of three key components: 1) It divides the teacher’s top items into two subsets; 2) It computes CE loss on items selected from both the teacher’s and the student’s top items; 3) An adaptive mechanism is proposed to combine the losses on these subsets. To validate the effectiveness of these key components, we design four variants: 1) RCE-KD w/o sep does not compute losses for $(\mathcal{Q}_{u}^{T})_{1}$ and $(\mathcal{Q}_{u}^{T})_{1}$, separately. It only computes CE loss on $\mathcal{A}^{u}\cup\mathcal{Q}_{u}^{S}$; 2) RCE-KD w/o S only aligns the predictions on the teacher’s predicted top items, i.e., $\mathcal{Q}_{u}^{T}$; 3) Similarly, RCE-KD w/o T only aligns the predictions on the student’s predicted top items, i.e., $\mathcal{Q}_{u}^{S}$; 4) RCE-KD w/ const replaces the adaptive weight derived from Eq.(10) with a constant hyperparameter $\gamma$. We also compare with CE loss computed on the full item set, denoted as full-CE, to validate the effectiveness of our method.
Figure 3 shows the results of these four variants on Gowalla and Yelp, and three Teacher/Student settings. The results of the remaining settings are provided in Appendix C.3. We find that all variants are inferior to the original RCE-KD, which demonstrates the effectiveness of all key components. Moreover, RCE-KD w/o S usually performs worse than RCE-KD w/o T. We believe that the reason is that the top items given by the student can exactly satisfy Assumption 4.3, while the top items given by the teacher do not. The superiority of RCE-KD w/ const over RCE-KD w/o T demonstrates the necessity of involving top items from both the student and the teacher. The superiority of RCE-KD over RCE-KD w/ const and RCE-KD w/o sep validates the effectiveness of our adaptive weighting scheme and the necessity of splitting out the two subsets and treating them separately. Finally, our method performs even slightly above full-CE in most cases, due to a tighter bound on NDCG than full-item CE (Xu et al., 2024a).